# SwiftKV: Fast Prefill-Optimized Inference with Knowledge-Preserving Model Transformation

### v2

@@ 3.4 Knowledge Recovery
While SingleInputKV preserves all the original parameters, it re-wires the architecture so that the KV cache projections may receive different inputs. We found that this re-wiring (and AcrossKV) requires fine-tuning to recover the original capabilities from the modified model. As we only change the computation of the attention part for layer $>l$, this can be achieved by fine-tune just the ${\mathbf{W}}_{QKV}$ weight matrices from the $(l+1)$-th layer onwards. However, instead of directly fine-tuning these parameters using standard LM loss, we find that distilling using the output logits of the original model allows for better knowledge recovery (see Sec. 5 for more details).
Implementing the Distillation. Since only a few ${\mathbf{W}}_{QKV}$ parameters need fine-tuning, we are able to do a memory efficient parameter-sharing based distillation. More specifically, we keep a single copy of the original model weights in memory that are frozen during training, and add an extra trainable copy of the ${\mathbf{W}}_{QKV}$ parameters for layers $>l$ initialized using the original model (See Fig. 1).
During the training, we create two forward modes for the later layers $>l$, one with original frozen parameters using original architecture, and another with the SwiftKV re-wiring using new QKV projections i.e.,
${\bf y}_{teacher}={\mathbf{M}}({\bf x},SwiftKV=False),~{}~{}\text{and}~{}~{}~{}{\bf y}_{student}={\mathbf{M}}({\bf x},SwiftKV=True),$ (3)
where ${\bf y}_{\cdot}$ is the final logits, ${\mathbf{M}}$ is the model, and ${\bf x}$ is the input. Afterwards, we apply the standard distillation loss ($L$) upon the outputs with temperature ($\tau$) using (Hinton et al., 2015). After the distillation, the original KV projection layers $>l$ are discarded during inference.
This method allows us to perform the distillation for Llama-3.1-8B-Instruct on 680M tokens of data in 3 hours using 8 H100 GPUs, and Llama-3.1-70B-Instruct in 5 hours using 32 H100 GPUs across 4 nodes.
@@ 3.5 Optimized Implementation for Inference
LLM serving systems can be complex and incorporate many simultaneous optimizations at multiple layers of the stack, such as PagedAttention (Kwon et al., 2023), Speculative Decoding (Leviathan et al., 2023), SplitFuse (Holmes et al., 2024; Agrawal et al., 2024), and more. One benefit of SwiftKV is that it makes minimal changes to the model architecture, limited to only a few linear projection layers. This means that SwiftKV can easily be integrated into existing serving systems without implementing new kernels (e.g. for custom attention operations or sparse computation) or novel inference procedures.
Implementation in vLLM. To realize the performance benefits of SwiftKV, we integrated it with vLLM . Our implementation is compatible with vLLM’s chunked prefill, which processes prefill tokens in chunks and may mix prefills and decodes in each minibatch. During each forward pass, after completing layer $l$, the KV-cache for the remaining layers ($>l$) are immediately computed, and only the decode tokens are propagated through the rest of the model layers.
GEMM and Memory Optimizations. Upon this basic implementation, we implemented two additional optimizations. First, SingleInputKV fusion: instead of computing the KV cache ${\mathbf{KV}}_{j}$ for each layer $j>l$ one at a time, we fused all ${\mathbf{W}}_{KV}^{j}$ into one large weight matrix ${\mathbf{W}}_{KV}^{j>l}$ so that their KV cache can be computed with a single efficient GEMM operation. Second, AcrossKV reduction: we modified vLLM to only allocate one layer’s KV-cache for each group of merged layers, which realizes the memory gains of AcrossKV.
@@ 4 Main Results
4.1 Setup
Training and Evaluation. We use Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct as our base models for SwiftKV. Our training datasets include a mixture of the full supervised training data from HuggingFaceH4/ultrachat_200k (Ding et al., 2023) and teknium/OpenHermes-2.5 (Teknium, 2023). We evaluated model quality using a modified LM-Eval-Harness (Gao et al., 2024) [^3] https://github.com/neuralmagic/lm-evaluation-harness/tree/llama_3.1_instruct due to its support for the custom prompt format of Llama-3.1, particularly for MMLU and MMLU-CoT (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), and Arc-Challenge (Clark et al., 2018). For more details, please see Appendix B.
@@ 4.1 Setup
Compression Metrics. For prefill computation, we report the approximate reduction as $(L-l)/L$ due to SingleInputKV, and for KV cache, we report the exact memory reduction due to AcrossKV. For example, SwiftKV with SingleInputKV ($l=L/2$) and 4-way AcrossKV is reported as 50% prefill compute reduction and 37.5% KV cache memory reduction. We further study how these theoretical compute and memory reductions translate into end-to-end inference improvements in Sec. 4.3.
Inference Performance. In our inference evaluation, we focus on two common scenarios: batch-inference for cost sensitive scenarios and interactive-inference for latency sensitive scenario.
Batch-Inference When processing text in bulk or when serving a model under high usage demand, it is important to achieve high combined throughput in terms of input + output tokens processed per second. For bulk processing, the combined throughput determines the time it takes to finish the job. For interactive use, it determines the volume of concurrent users and requests that can be supported per unit of hardware. In both scenarios, the combined throughput is a key determinant of the cost of serving the model.
Interactive-Inference In interactive scenarios (e.g., chatbots, copilots), not only the combined throughput is important, but also metrics that define the end-user experience. Chief upon them are the time-to-first-token (TTFT) and time-per-output-token (TPOT). TTFT is the time between the user sending a message and receiving the first token in the response. TPOT is the time between each output token after the first token has been received. Low TTFT and TPOT are desirable by interactive applications to deliver smooth usage experiences.
For all experiments on Llama-3.1-8B-Instruct, we use 1 NVIDIA H100 GPU with 80GB of memory, and for all experiments on Llama-3.1-70B-Instruct, we use 4 NVIDIA H100 GPUs running the model with 4-way tensor parallelism. We provide the full hardware and vLLM configurations in Appendix B.2
4.2 Model Quality with Prefill Compute Reduction
Llama-3.1-8B-Instruct. The top rows of  Table 2 show that SwiftKV can preserve model quality well until 50% prefill reduction using SingleInputKV. For 25% prefill reduction, the accuracy degradation is only about 0.12 points and for 50% reduction, the gap is about 1 point  [^4] Note that we did not try to find the best training recipe, regarding to either training data (e.g., we did not include any math or coding datasets) or training pipeline (e.g., we did not include reinforce-learning rated steps, like DPO/RLHF). Yet, the quality of SwiftKV is close to the original base Llama-3.1-8/70B-Instruct models. In Sec. D, we show that better data recipe could boost the model performance and close the quality gap further.. When we push to 62.5% reduction (i.e. SingleInputKV with $l=12$ and $L=32$), the accuracy drops to 66.09 points, which is significantly lower than the baseline. This can be explained by the drop in activation similarity from 0.61 to 0.51 between layer 16 to layer 12 (Fig. 2).
The bottom rows of Table 2 show the model quality when adding AcrossKV to 50% SingleInputKV. From pure SingleInputKV to 2-way AcrossKV, the accuracy drops about 0.9 points with 25% KV cache reduction. The accuracy drops by another 0.32, going from 2-way to 8-way sharing, and 0.62 when going all the way to 16-way sharing. Particularly, for the extreme case, i.e., using a single KV cache for all remaining layers, the accuracy is only about 2.5 points lower than pure SingleInputKV, and could be useful for more memory constrained cases, e.g., embedding and/or mobile devices.
Furthermore, the design of AcrossKV is complementary to many existing KV cache compression methods. In Sec. 5.3, we show that AcrossKV can be combined with quantization to achieve 62.5% reduction in KV cache memory with only a 1-point accuracy gap compared to SingleInputKV only.
Table 2: Llama-3.1-8B-Instruct
Model SingleInputKV AcrossKV Arc-Challenge Winogrande HelloSwag truthfulqa MMLU MMLU-CoT GSM-8K Avg. (Prefill Reduction) (Cache Reduction) 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots Baseline N/A N/A 82.00 77.90 80.40 54.56 67.90 70.63 82.56 73.71 SwiftKV ✓(25%) 82.08 77.98 80.63 54.59 67.95 70.45 81.43 73.59 SwiftKV ✓(50%) 80.38 78.22 79.30 54.54 67.30 69.73 79.45 72.70 SwiftKV ✓(62.5%) 71.76 75.77 78.21 52.73 61.55 53.68 68.92 66.09 SwiftKV ✓(50%) 2-way (25%) 80.29 77.82 79.03 54.66 66.96 68.39 75.59 71.82 SwiftKV ✓(50%) 4-way (37.5%) 79.35 77.51 78.44 54.96 65.71 67.75 76.72 71.49 SwiftKV ✓(50%) 8-way (43.75%) 79.18 77.19 77.38 54.79 65.73 66.88 72.33 70.50 SwiftKV ✓(50%) 16-way (46.875%) 78.24 76.80 76.87 56.86 64.65 65.86 72.25 70.22
Table 3: Llama-3.1-70B-Instruct
Model SingleInputKV AcrossKV Arc-Challenge Winogrande HelloSwag truthfulqa MMLU MMLU-CoT GSM-8K Avg. Prefill Reduction Cache Reduction 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots Baseline N/A N/A 93.34 85.16 86.42 59.95 83.97 86.21 95.15 84.31 SwiftKV ✓(25%) 93.00 84.69 85.98 59.43 82.82 85.81 95.07 83.83 SwiftKV ✓(50%) 93.09 83.82 84.45 58.40 82.51 85.00 93.56 82.98 SwiftKV ✓(50%) 2-way (25%) 92.92 82.95 84.10 57.79 82.66 84.55 93.48 82.63 SwiftKV ✓(50%) 4-way (37.5%) 92.92 83.74 84.72 58.28 82.60 84.79 93.71 82.96
Llama-3.1-70B-Instruct.  Table 3 shows that with 50% prefill reduction using SingleInputKV, Llama-3.1-70B-Instruct incurs a 1.3 point drop in accuracy which is slightly higher than the results of Llama-3.1-8B-Instruct. However, Llama-3.1-70B-Instruct is more resilient to AcrossKV, incurring less than a 0.35 point drop in accuracy even for 4-way sharing across layers.
Table 4: Summary of all other models we evaluated, only average scores shown.
@@ 4.2 Model Quality with Prefill Compute Reduction
Model SingleInputKV AcrossKV Llama-3.2-3B- Llama-3.1-405B- Mistral-Small- Deepseek-V2- Prefill Reduction Cache Reduction Instruct Instruct (FP8) Instruct-2409 Lite-Chat Baseline N/A N/A 66.47 86.6 78.2 64.12 SwiftKV ✓(~25%) 66.55 64.19 SwiftKV ✓(40–50%) 65.55 85.9 77.2 63.51 SwiftKV ✓(40–50%) 2-way (25%) 65.13 63.07 SwiftKV ✓(40–50%) 4-way (37.5%) 65.19 76.7 59.32
Other models and diverse architectures. In addition to Llama-3.1-Instruct 8B and 70B, we integrated and evaluated four other models with SwiftKV, which span a more diverse spectrum of model architectures. Of particular note, Llama-3.1-405B-Instruct is run at 8-bit precision using W8A8 quantization, and Deepseek-V2-Lite-Chat is an mixture-of-experts model that also implements a novel attention mechanism that compresses its keys and values using a latent vector (DeepSeek-AI et al., 2024).
Table 4 summarizes the results, and the full per-task evaluation scores can be found in Appendix B.3. Overall, we find that SwiftKV generalizes well to all these models, with few minor exceptions. First, Llama-3.2-3B-Instruct experiences a steeper quality degradation at 50% SingleInputKV, but performs well at 40% SingleInputKV (detailed comparison in Appendix B.3). Second, Deepseek-V2-Lite-Chat experiences a steeper quality degradation at 4-way AcrossKV, but performs well at 2-way AcrossKV.

### v3

@@ 3.4 Knowledge Recovery
While SwiftKV preserves all the original parameters, it re-wires the architecture so that the KV cache projections may receive different inputs. We found that this re-wiring (and AcrossKV) requires fine-tuning to recover the original capabilities from the modified model. Since we only change the KV projections for layer $>l$, this can be achieved by fine-tuning just the ${\mathbf{W}}_{QKV}$ weight matrices from the $(l+1)$-th layer onwards. However, instead of directly fine-tuning these parameters using standard LM loss, we find that distilling using the output logits of the original model allows for better knowledge recovery (see Sec. 5 for more details).
Additionally, we found that limiting the training to just ${\mathbf{W}}_{QKV}$ achieves better accuracy, which aligns with prior hypotheses that LLM knowledge is primarily stored in their MLP layers Meng et al. (2024); Geva et al. (2021); Elhage et al. (2021). We further explore this in Sec. 5.2. An added benefit is that these parameters are typically <10% of the total for popular GQA models (e.g., Llama, Mistral, Qwen), allowing for very efficient distillation.
Efficient Distillation.
Since only a few ${\mathbf{W}}_{QKV}$ parameters need training, we can keep just a single copy of the original model weights in memory that are frozen during training, and add an extra trainable copy of the ${\mathbf{W}}_{QKV}$ parameters for layers $>l$ initialized using the original model (See Fig. 1).
During training, we create two modes for the later layers $>l$, one with original frozen parameters using original architecture, and another with the SwiftKV re-wiring using new QKV projections i.e.,
${\bf y}_{teacher}$ $={\mathbf{M}}({\bf x},SwiftKV=False),$ (3)
${\bf y}_{student}$ $={\mathbf{M}}({\bf x},SwiftKV=True),$
where ${\bf y}_{\cdot}$ is the final logits, ${\mathbf{M}}$ is the model, and ${\bf x}$ is the input. Afterwards, we apply the standard distillation loss Hinton et al. (2015) on the outputs. After the distillation, the original KV projection layers $>l$ are discarded during inference.
This method allows us to distill Llama-3.1-8B-Instruct on 680M tokens of data in 3 hours using 8 H100 GPUs, and Llama-3.1-70B-Instruct in 5 hours using 32 H100 GPUs across 4 nodes. In contrast, many prune-and-distill Sreenivas et al. (2024) and layer-skipping Elhoushi et al. (2024) methods require much larger datasets (e.g. 10–100B tokens) and incur greater accuracy gaps than SwiftKV.
@@ 3.5 Optimized Implementation for Inference
LLM serving systems can be complex and incorporate many simultaneous optimizations at multiple layers of the stack, such as PagedAttention (Kwon et al., 2023), Speculative Decoding (Leviathan et al., 2023), SplitFuse (Holmes et al., 2024; Agrawal et al., 2024), and more. A benefit of SwiftKV is that it makes minimal changes to the model architecture, so it can be integrated into existing serving systems without implementing new kernels (e.g. for custom attention operations or sparse computation) or novel inference procedures.
Implementation in vLLM and SGLang.
To show that the theoretical compute reductions of SwiftKV translates to real-world savings, we integrated it with vLLM Kwon et al. (2023) and SGLang Zheng et al. (2024). Our implementation is compatible with chunked prefill Holmes et al. (2024); Agrawal et al. (2024), which mixes chunks of prefill tokens and decode tokens in each minibatch. During each forward pass, after completing layer $l$, the KV-cache for the remaining layers ($>l$) are immediately computed, and only the decode tokens are propagated through the rest of the model layers.
@@ 4 Main Results
We evaluated SwiftKV in terms of model accuracy (Sec. 4.1) compared to the original model and several baselines, and end-to-end inference performance (Sec. 4.2) in a real serving system.
Distillation datasets.
Our dataset is a mixture of Ultrachat (Ding et al., 2023), SlimOrca (Lian et al., 2023), and OpenHermes-2.5 (Teknium, 2023), totaling roughly 680M Llama-3.1 tokens. For more details, please see Appendix A.1.
SwiftKV Notation.
@@
For prefill computation, we report the approximate reduction as $(L-l)/L$ due to SwiftKV, and for KV cache, we report the exact memory reduction due to AcrossKV. For example, SwiftKV ($l=L/2$) and 4-way AcrossKV is reported as 50% prefill compute reduction and 37.5% KV cache memory reduction.
4.1 Model Quality Impact of SwiftKV
Table 2: All SwiftKV model quality evaluations. For FFN-SkipLLM, we set the candidate layers to be skipped to be from 35–8% depth in each model, which reflects the settings in their paper. The prefill reduction % represents just the fraction of MLP layer skipped, and varies between models and tasks since it is adaptively determined during inference.
@@ 4.1 Model Quality Impact of SwiftKV
Model SwiftKV AcrossKV Arc-Challenge Winogrande Hellaswag TruthfulQA MMLU MMLU-CoT GSM8K-CoT Avg. (Prefill Reduction) (Cache Reduction) 0-shot 5-shot 10-shot 0-shot 5-shot 0-shot 8-shot Llama-3.1-8B-Instruct Baseline 82.00 77.90 80.40 54.56 67.90 70.63 82.56 73.71 SwiftKV ✓(25%) 82.08 77.98 80.63 54.59 67.95 70.45 81.43 73.59 SwiftKV ✓(50%) 80.38 78.22 79.30 54.54 67.30 69.73 79.45 72.70 SwiftKV ✓(62.5%) 71.76 75.77 78.21 52.73 61.55 53.68 68.92 66.09 SwiftKV ✓(50%) 2-way (25%) 80.29 77.82 79.03 54.66 66.96 68.39 75.59 71.82 SwiftKV ✓(50%) 4-way (37.5%) 79.35 77.51 78.44 54.96 65.71 67.75 76.72 71.49 SwiftKV ✓(50%) 8-way (43.75%) 79.18 77.19 77.38 54.79 65.73 66.88 72.33 70.50 SwiftKV ✓(50%) 16-way (46.875%) 78.24 76.80 76.87 56.86 64.65 65.86 72.25 70.22 FFN-SkipLLM (12-19%) 81.4 74.11 73.94 54.55 67.65 64.12 78.62 70.62 Llama-3.1-70B-Instruct Baseline 93.34 85.16 86.42 59.95 83.97 86.21 95.15 84.31 SwiftKV ✓(25%) 93.00 84.69 85.98 59.43 82.82 85.81 95.07 83.83 SwiftKV ✓(50%) 93.09 83.82 84.45 58.40 82.51 85.00 93.56 82.98 SwiftKV ✓(50%) 2-way (25%) 92.92 82.95 84.10 57.79 82.66 84.55 93.48 82.63 SwiftKV ✓(50%) 4-way (37.5%) 92.92 83.74 84.72 58.28 82.60 84.79 93.71 82.96 Nemotron-51B (28%) (50%) 91.47 84.45 85.68 59.02 81.74 83.86 93.25 82.78 Llama-3.1-405B-Instruct (FP8) Baseline 94.7 87.0 88.3 64.7 87.5 88.1 96.1 86.6 SwiftKV ✓(50%) 94.0 86.3 88.1 64.2 85.7 87.5 95.2 85.9 Llama-3.2-3B-Instruct Baseline 75.17 68.59 73.32 51.45 62.01 62.48 72.32 66.47 SwiftKV ✓(25%) 75.59 69.77 72.34 52.80 61.89 62.39 71.11 66.55 SwiftKV ✓(40%) 75.34 68.98 71.37 51.10 61.80 61.62 68.68 65.55 SwiftKV ✓(50%) 71.25 68.75 70.77 51.29 59.63 59.94 67.02 64.09 SwiftKV ✓(40%) 2-way (25%) 74.82 68.66 71.41 50.67 61.55 61.03 67.77 65.13 SwiftKV ✓(40%) 4-way (37.5%) 75.59 69.21 70.79 50.89 61.35 60.82 67.70 65.19 FFN-SkipLLM (8-16%) 74.57 66.38 67.55 49.57 60.95 61.24 69.75 64.28 Mistral-Small-Instruct-2409 Baseline 84.12 84.68 87.27 56.85 73.33 74.86 86.50 78.23 SwiftKV ✓(25%) 84.04 84.84 87.03 55.97 72.88 74.69 85.21 77.80 SwiftKV ✓(50%) 83.53 83.97 86.30 55.63 72.91 74.04 84.30 77.24 SwiftKV ✓(50%) 2-way (25%) 83.36 84.05 86.22 56.20 72.30 73.70 84.68 77.21 SwiftKV ✓(50%) 4-way (37.5%) 82.93 83.82 86.17 56.00 72.29 73.00 82.48 76.66 FFN-SkipLLM (34-37%) 65.61 72.61 59.80 53.52 64.20 2.16 2.12 45.71 Deepseek-V2-Lite-Chat Baseline 65.53 74.66 81.56 50.98 56.86 50.61 68.69 64.12 SwiftKV ✓(25%) 65.44 75.05 81.52 50.53 56.91 50.92 68.99 64.19 SwiftKV ✓(45%) 65.61 73.95 80.82 50.20 56.33 51.56 66.11 63.51 SwiftKV ✓(45%) 2-way (25%) 65.52 74.26 80.23 49.85 55.59 50.51 65.57 63.07 SwiftKV ✓(45%) 4-way (37.5%) 61.34 75.21 79.80 48.39 54.82 30.80 64.89 59.32 FFN-SkipLLM (30-32%) 10.49 58.41 49.34 50.69 4.56 0.01 0.30 24.83 Qwen2.5-14B-Instruct Baseline 62.29 79.32 85.04 69.07 76.58 79.04 90.37 77.38 SwiftKV ✓(25%) 62.03 79.00 84.63 68.39 76.09 78.64 84.83 76.23 SwiftKV ✓(50%) 56.91 77.26 82.71 60.76 64.40 68.20 79.30 69.93 SwiftKV ✓(25%) 2-way (25%) 61.43 79.71 85.22 69.33 76.25 78.88 84.23 76.43 SwiftKV ✓(25%) 4-way (37.5%) 59.13 80.89 84.92 68.75 75.70 78.84 82.78 75.85 FFN-SkipLLM (7–21%) 53.24 73.09 65.10 59.78 73.55 62.22 50.79 62.53 DarwinLM-8.4B (40%) 49.32 70.96 74.95 41.99 12.46 0.00 1.90 35.94
Table 2 shows the quality results of all models we evaluated, including Llama-3.1-Instruct, Qwen2.5-14B-Instruct, Mistral-Small, and Deepseek-V2. Of these models, we note that the Llama models span two orders of magnitude in size (3B to 405B), Llama-3.1-405B-Instruct uses FP8 (W8A16) quantization, and Deepseek-V2-Lite-Chat is a mixture-of-experts model that implements a novel latent attention mechanism (DeepSeek-AI et al., 2024).
We also compare with three baselines: (1) FFN-SkipLLM (Jaiswal et al., 2024), a training-free method for skipping FFN layers (no attention layers are skipped) based on hidden state similarity, (2) Llama-3.1-Nemotron-51B-Instruct Sreenivas et al. (2024), which is pruned and distilled from Llama-3.1-70B-Instruct using neural architecture search on 40B tokens, and (3) DarwinLM-8.4B Tang et al. (2025), which is pruned and distilled from Qwen2.5-14B-Instruct using 10B tokens.
SwiftKV.
For Llama, Mistral, and Deepseek, we find the accuracy degradation for 25% SwiftKV is less than 0.5% from the original models (averaged across tasks). Additionally, the accuracy gap is within 1–2% even at 40–50% SwiftKV. Beyond 50% SwiftKV, model quality drops quickly. For example, Llama-3.1-8B-Instruct incurs a 7% accuracy gap at 62.5% SwiftKV. We find that Qwen suffers larger degradations, at 1.1% for 25% SwiftKV and 7.4% for 50% SwiftKV, which may be due to Qwen models having lower simularity between layer at 50–75% depth (Fig. 2). Even still, SwiftKV performs much better than FFN-SkipLLM and DarwinLM-8.4B, which suffer massive 15% and 42% drops from the baseline model, respectively.
AcrossKV.
The accuracy impact of AcrossKV is also minimal. Starting from 25–50% SwiftKV, adding 2-way AcrossKV (20–25% KV cache reduction) further degrades average task accuracy by at most 1% across all models. Pushing to 4-way AcrossKV, Deepseek-V2-Lite-Chat experiences a steep accuracy drop from 63.07% to 59.32%, while other models experience smaller drops. Notably, we found that Llama-3.1-8B-Instruct still achieves 70.22% average accuracy at 16-way AcrossKV, meaning all the last half of layers share a single layer of KV cache. Furthermore, the design of AcrossKV is complementary to many existing KV cache compression methods. In Sec. B.1, we show that AcrossKV can be combined with quantization to achieve 62.5% reduction in KV cache memory.
SwiftKV vs Baselines.
SwiftKV outperforms FFN-SkipLLM across all scenarios we tested. FFN-SkipLLM skips only MLPs for prefill and decode tokens, while SwiftKV skips both MLP and attention layers for prefill tokens. Still, FFN-SkipLLM sees large degradations for Mistral, Deepseek, and Qwen, even at 7–37% of MLPs skipped. For Llama models, skipping under 20% of of the MLP layers using FFN-SkipLLM still underperforms SwiftKV skipping 50% of MLP and attention layers.
Compared with Nemotron-51B and DarwinLM-8.4B, 50% SwiftKV reduces more prefill while achieving higher accuracies. Also, Nemotron-51B is distilled on 40B tokens, and DarwinLM-8.4B on 10B tokens, while SwiftKV is distilled on <1B tokens and on <10% of the model parameters. When prefill compute is substantial (e.g., many enterprise applications), SwiftKV is the clear choice for reducing cost without sacrificing accuracy.